Solve the system of linear equations by the elimination method.
If 3/11x - 1/2y = 5/11
and 2/5x - 1/3y = 2/15
please provide the steps to this problem I am completely lost

Answer by stanbon(75887)

Solve the system of linear equations by the elimination method.
3/11x - 1/2y = 5/11
2/5x - 1/3y = 2/15
-------------------------
Multiply thru 1st by 22
Multiply thru 2nd by 15
-------------
6x - 11y = 10
6x - 5y = 2
------
Subtract the top eq. from the bottom and solve for "y":
6y = -8
y = -4/3
-------------------
Substitute for "y" in 6x-5y=2 and solve for "x":
6x -5(-4/3) = 2
6x + 20/3 = 6/3
6x = -14/3
x = -14/18 = -7/9
====================
Cheers,
Stan H.